OpenStudy (anonymous):

f(x)=10x-9 g(x)=-16x^2+x-11 find g(f(x))

hero (hero):

f(x) = 10x-9 g(x) = -16x^2 + x - 11 g(f(x)) = g(10x-9) = -16(10x-9)^2 + (10x-9) - 11 = -16(100x^2-180x + 81) + 10x - 9 - 11 = -1600x^2 + 2880x - 1296 + 10x - 20 = -1600x^2 + 2890x - 1316 Whew!

hero (hero):

So g(f(x)) = -1600x^2 + 2890x - 1316

hero (hero):

It's a little messy this time, but same concept

OpenStudy (anonymous):

you just plug f(x)=10x-9 into the g(x) and combine?

hero (hero):

Yes

OpenStudy (anonymous):

THANK YOU SO MUCH

hero (hero):

You're a good student. If only others could understand me so quickly. Usually, I would have to go on a long tedious journey (without knowing whether or not I'll reach the desired destination) to explain things. But you are clearly a quick learner.

OpenStudy (anonymous):

hahah thanks (: i'm a visual/hands-on learner so once i see someone do it and practice once i'm usually good (: thank you so much ! you just saved a section of my exam grade. hahah

hero (hero):

:D Glad to assist